# Taming Throughput-Latency Tradeoff in LLM Inference with Sarathi-Serve

### v1

@@ 3.1 Throughput-Latency Trade-off
Figure 4: A generation stall occurs when one or more prefills are scheduled in between consecutive decode iterations of a request. A, B, C and D represent different requests. Subscript $d$ represents a decode iteration, $p$ represents a full prefill and $p0$, $p1$ represent two chunked prefills of a given prompt. vLLM induces generation stalls by scheduling as many prefills as possible before resuming ongoing decodes. Despite supporting hybrid batches, Orca cannot mitigate generation stalls because the execution time of batches containing long prompts remains high. FasterTransformer is free of generation stalls as it finishes all ongoing decodes before scheduling a new prefill but compromises on throughput due to low batch size in many decode iterations. In contrast, Sarathi-Serve generates a schedule that eliminates generation stalls yet delivers high throughput.
Iteration-level batching improves system throughput but we show that it comes at the cost of high TBT latency due to a phenomenon we call as generation stalls as discussed below.
Figure 4 compares different scheduling policies. The example shows a timeline (left to right) of requests A, B, C and D. Requests A and B are in decode phase at the start of the interval and after one iteration, requests C and D enter the system. Orca and vLLM both use FCFS iteration-level batching with eager admission of prefill requests (lines 8-9 in Algorithm 2) but differ in their batch composition policy. Orca supports hybrid batches composed of both prefill and decode requests whereas vLLM only supports batches that contain either all prefill or all decode requests. Irrespective of this difference, both Orca and vLLM can improve throughput by maximizing the batch size in subsequent decode iterations. However, eagerly scheduling prefills of requests C and D delays the decodes of already running requests A and B because an iteration that computes one or more prefills can take several seconds depending on the lengths of input prompts. Therefore, prefill-prioritizing schedulers can introduce generation stalls for ongoing decodes resulting in latency spikes due to high TBT.
@@
One way to reduce latency spikes in iteration-level batching systems is to use smaller batch sizes as recommended in Orca [71]. However, lowering batch size adversely impacts throughput as shown in Figure 3 and therefore existing systems are forced to trade throughput or latency depending on the desired SLOs.
@@
Takeaway-2: The interleaving of prefills and decodes involves a trade-off between throughput and latency for current LLM inference schedulers. State-of-the-art systems today use prefill-prioritizing schedules that trade TBT latency for high throughput.
3.2 Low Compute Utilization during Decodes
Low compute utilization during the decode phase is a waste of GPU’s processing capacity. To understand this further, we analyze the arithmetic intensity of prefill and decode iterations. Figure 5 shows that the majority of the time in LLM inference is spent in linear operators. Therefore we focus our analysis on linear operators of the FFN module.
Figure 6: Arithmetic intensity trend for LLaMA2-70B linear operations with different number of token running on four A100s. Decode batches have low arithmetic intensity i.e., they are bottlenecked by memory fetch time, leading to low compute utilization. Prefill batches are compute bound with sub-optimal bandwidth utilization. Sarathi forms hybrid batches by combining decodes and prefill chunks to maximize both compute and bandwidth utilization.
Matrix multiplication kernels overlap memory accesses along with computation of math operations. The total execution time of an operation can be approximated to $T=\text{max}(T_{\text{math}},T_{\text{mem}})$, where $T_{\text{math}}$ and $T_{\text{mem}}$ represent the time spent on math and memory fetch operations respectively. An operation is considered memory-bound if $T_{\text{math}}<T_{\text{mem}}$. Memory-bound operations have Model FLOPs Utilization (MFU) [32], on the other hand, compute-bound operations have low Model Bandwidth Utilization (MBU). When $T_{\text{math}}=T_{\text{mem}}$, both compute and memory bandwidth utilization are maximized. Arithmetic intensity quantifies the number of math operations performed per byte of data fetched from the memory. At the optimal point, the arithmetic intensity of operation matches the FLOPS-to-Bandwidth ratio of the device. Figure 6 shows arithmetic intensity as a function of the number of tokens in the batch for linear layers in LLaMA2-70B running on four A100 GPUs. Prefill batches amortize the cost of fetching weights of the linear operators from HBM memory to GPU cache over a large number of tokens, allowing it to have high arithmetic intensity. In contrast, decode batches have very low computation intensity. Figure 7 shows the total execution time of linear operators in an iteration for LLaMA2-70B as a function of the number of tokens. Note that execution time increases only marginally in the beginning i.e., as long as the batch is in a memory-bound regime, but linearly afterwards i.e., when the batch becomes compute-bound. [^2] Theoretically, we expect the operators to become compute-bound at $\sim$200 tokens on A100 GPUs, however, in practice we observe that it happens at $\sim$500-600 tokens for higher tensor parallel dimensions due to fixed overheads.
Takeaway-3: Decode batches operate in memory-bound regime leaving compute underutilized. This implies that more tokens can be processed along with a decode batch without significantly increasing the latency of the decode iteration.
Figure 7: Linear layer execution time as function of number of tokens in a batch for LLaMA2-70B on A100(s) with different tensor parallel degrees. When the number of tokens is small, the the execution time is dictated by the cost of fetching weights from HBM memory. As a result, the execution time is largely stagnant in the 128-512 tokens range, especially for higher tensor parallel degrees. Once the number of tokens in the batch cross a critical threshold, the operation become compute bound and the runtime increases linearly with number of tokens.
@@ 4 Sarathi-Serve: Design and Implementation
We now discuss the design and implementation of Sarathi-Serve which uses the techniques chunked-prefills defined in our prior work Sarathi [27] to create a stall-free batching scheduler optimized for online inference serving.
@@ 4.1 Chunked-prefills
In our previous work Sarathi, we presented a technique called chunked-prefills which allows computing large prefills in small chunks across several iterations. In Sarathi-Serve, we leverage this mechanism to form batches with appropriate number of tokens such that we can utilize the compute potential in decode batches without violating the TBT SLO.
@@
2:Initialize token_budget, $\tau$leftarrow compute_token_buget($T_{max}$)
3:Initialize batch_num_tokens, $n_{t}$leftarrow 0
@@
7:      if is_prefill_complete($R$then
@@
10:      if not is_prefill_complete($R$then
@@ 4.2 Stall-free batching
The Sarathi-Serve scheduler is an iteration-level scheduler that leverages chunked-prefills and coalescing of prefills and decodes from Sarathi to improve throughput while minimizing latency.
Unlike Orca and vLLM which stall existing decodes to execute prefills, Sarathi-Serve leverages the arithmetic intensity slack in decode iterations to execute prefills without delaying the execution of decode requests in the system. We call this approach stall-free batching (Algorithm 3). Sarathi-Serve first calculates the budget of maximum number of tokens that can be executed in a batch based on user specified SLO. We describe the considerations involved in determining this token budget in depth in §4.3. In every scheduling iteration, we first pack all the running decodes in the next batch (lines 6-8 in Algorithm 3). After that, we include any partially completed prefill (lines 9-12). Only after all the running requests have been accommodated, we admit new requests (lines 13-20). When adding prefill requests to the batch, we compute the maximum chunk size that can be accommodated within the leftover token budget for that batch (lines 11, 15). By restricting the computational load in every iteration, stall-free batching ensures that decodes never experience a generations stall due to a co-running prefill chunk. We compare the latency for hybrid batches with and without chunked prefills in Figure 8. Naive hybrid batching leads to dramatic increase of up to 28.3$\times$ in the TBT latency compared to a decode-only batch. In contrast, Sarathi-Serve provides a much tighter bound on latency with chunking.
Figure 4 shows the scheduling policy of Sarathi-Serve in action, for the same example used in §3.1. The first iteration is decode-only as there are no prefills to be computed. However, after a new request C enters the system, Sarathi-Serve first splits the prefill of C into two chunks and schedules them in subsequent iterations. At the same time, with stall-free batching, it coalesces the chunked prefills with ongoing decodes of A and B. This way, Sarathi-Serve stalls neither decodes nor prefills unlike existing systems, allowing Sarathi-Serve to be largely free of latency spikes in TBT without compromising throughput.
@@ 4.3 Determining Token Budget
During the computation of chunked-prefills, the attention operation for every chunk of a prompt needs to access the KV-cache of all prior chunks of the same prompt. This results in increased memory reads from the GPU HBM even though the computational cost is unchanged. For example, if a prefill sequence is split into $N$ chunks, then the first chunk’s KV-cache is loaded $N-1$ times, the second chunk’s KV-cache is loaded $N-2$ times, and so on. However, the overhead due to increased attention time is not significant as attention computation is a relatively small fraction of the overall time. We present a detailed study of the overheads of chunked-prefills in §5.2.
@@
Another factor that influences the choice of token budget is the tile-quantization effect [13]. GPUs compute matmuls by partitioning the given matrices into tiles and assigning them to different thread blocks for parallel computation. Here, each thread block refers to a group of GPU threads and computes the same number of arithmetic operations. Therefore, matmuls achieve maximum GPU utilization when the matrix dimensions are divisible by the tile size. Otherwise, due to tile-quantization, some thread blocks perform extraneous (wasted) computation [13]. We observe that tile-quantization can dramatically increase prefill computation time e.g., in some cases, using chunk size of 257 can increase prefill time by 32% compared to that with chunk size 256.
Therefore, selecting a suitable token budget is a two-fold decision. First, pick a token budget based on the desired TBT SLO. Next, ensure that the total number of tokens in a batch is a multiple of the tile size. The latter is relatively straightforward since the size of tile dimensions on GPUs is almost always a power of two.
@@ 5 Evaluation
    What is the maximum load a model replica can serve under various Service Level Objective (SLO) constraints with different inference serving systems?
@@
    What is the overhead of chunked-prefills?
@@
    What is the effect of each of chunked-prefills and stall-free batching in isolation as opposed to using them in tandem?
Implementation: We implement Sarathi-Serve on top of the open-source implementation of vLLM [50, 21]. The framework leverages PyTorch [54] and xFormers [22] for matrix multiplication and attention kernels respectively. We extend the base vLLM codebase to support various scheduling policies, chunked prefills, pipeline parallelism and an extensive telemetry system. We use NCCL [15] for both pipeline and tensor parallel communication.
@@ 5.1 Capacity Evaluation
We evaluate Sarathi-Serve, Orca and vLLM on all four models and both datasets under two different latency configurations: relaxed and strict. Similar to Patel et al. [55], to account for the intrinsic performance limitations of a model and hardware pair, we define the SLO on P99 TBT to be equal to 5$\times$ and 25$\times$ the execution time of a decode iteration for a single request (with prefill length of 4k and 32 batch size) running without any contention for the strict and relaxed settings, respectively. Table 3 shows a summary of the absolute SLO thresholds. Note that the strict target represents the target SLOs desired for interactive applications like chatbots. On the other hand, the relaxed configuration is exemplary of systems where the complete sequence of output tokens should be generated within a predictable time limit but the TBT constraints on individual tokens is not very strict. For all load experiments, we ensure that the maximum load is sustainable, i.e., the queuing delay does not blow up (we use a limit of 2 seconds median scheduling delay).
Figure 9 and Figure 10 show the results of our capacity experiments. Sarathi-Serve consistently outperforms Orca and vLLM in all cases across models and workloads. Under strict SLO, Sarathi-Serve can sustain up to 3.2×3.2$\times$3.2 × higher load compared to Orca (Mistral-7B, openchat_sharegpt4) and 2.7×2.7$\times$2.7 × higher load than vLLM under relaxed SLO (Yi-34B, openchat_sharegpt4). The serving capacity of the bigger 70B and 180B models drop significantly compared to smaller models. However, even for these bigger models, Sarathi-Serve consistently outperforms Orca and vLLM.
@@
We observe that in most scenarios, Orca and vLLM violate the P99 TBT latency SLO before they can reach their maximum serviceable throughput. Thus, we observe relaxing the latency target leads to considerable increase in their model serving capacity. In Sarathi-Serve, one can adjust the chunk size based on the desired SLO. Therefore, we use a strict token budget and split prompts into smaller chunks when operating under strict latency SLO. This reduces system efficiency marginally but allows us to achieve lower tail latency. On the other hand, when the latency constraint is relaxed, we increase the token budget to allow more efficient prefills. We use token budget of 2048 and 512 for all models under the relaxed and strict settings, respectively, except for the LLaMA2-70B relaxed configuration where we use token budget of 1536 to reduce the impact of pipeline bubbles. The system performance can be further enhanced by dynamically varying the token budget based on workload characteristics. We leave this exploration for a future work.
@@
We further notice that vLLM significantly outperforms Orca. The reason for this is two-fold. First, Orca batches prompts for multiple requests together (max sequence length * batch size compared to max sequence length in vLLM), which can lead to even higher tail latency in some cases. Second, vLLM supports a much larger batch size compared to Orca. The lower batch size in Orca is due to the lack of PagedAttention and the large activation memory footprint associated with processing batches with excessively large number of tokens.
@@
Note that the capacity of each system is higher for openchat_sharegpt4 dataset compared to the arxiv_summarization dataset. This is expected because prompts in the arxiv_summarization datasets are much longer - 7059 vs 1730 median tokens as shown in Table 2. The higher prompts makes Orca and vLLM more susceptible to latency violations due to higher processing time of these longer prefills.

### v3

@@ 3.2 Throughput-Latency Trade-off
Iteration-level batching improves system throughput but we show that it comes at the cost of high TBT latency due to a phenomenon we call generation stalls.
Figure 7 compares different scheduling policies. The example shows a timeline (left to right) of requests A, B, C and D. Requests A and B are in decode phase at the start of the interval and after one iteration, requests C and D enter the system. Orca and vLLM both use FCFS iteration-level batching with eager admission of prefill requests but differ in their batch composition policy. Orca supports hybrid batches composed of both prefill and decode requests whereas vLLM only supports batches that contain either all prefill or all decode requests. Irrespective of this difference, both Orca and vLLM can improve throughput by maximizing the batch size in subsequent decode iterations. However, eagerly scheduling prefills of requests C and D delays the decodes of already running requests A and B because an iteration that computes one or more prefills can take several seconds depending on the lengths of input prompts. Therefore, prefill-prioritizing schedulers can introduce generation stalls for ongoing decodes resulting in latency spikes caused by high TBT.
@@
One way to reduce latency spikes in iteration-level batching systems is to use smaller batch sizes as recommended in Orca [75]. However, lowering batch size adversely impacts throughput as shown in §2.2. Therefore, existing systems are forced to trade-off between throughput and latency depending on the desired SLOs.
@@
Takeaway-3: The interleaving of prefills and decodes involves a trade-off between throughput and latency for current LLM inference schedulers. State-of-the-art systems today use prefill-prioritizing schedules that trade TBT latency for high throughput.
Figure 8: A 2-way pipeline parallel iteration-level schedule in Orca across 4 requests (A,B,C,D) shows the existence of pipeline bubbles due to non-uniform batch execution times. Sarathi-Serve is able to minimize these stalls by creating uniform-compute batches.
3.3 Pipeline Bubbles waste GPU Cycles
Pipeline-parallelism (PP) is a popular strategy for cross-node deployment of large models, owing to its lower communication overheads compared to Tensor Parallelism (TP). A challenge with PP, however, is that it introduces pipeline bubbles or periods of GPU inactivity as subsequent pipeline stages have to wait for the completion of the corresponding micro-batch in the prior stages. Pipeline bubbles is a known problem in training jobs, where they arise between the forward and backward passes due to prior stages needing to wait for the backward pass to arrive. Micro-batching is a common technique used in PP training jobs to mitigate pipeline bubbles [33, 55, 49].
Inference jobs only require forward computation and therefore one might expect that micro-batching can eliminate pipeline bubbles during inference. In fact, prior work on transformer inference, such as, FasterTransformer [7] and FastServe [72] use micro-batches but do not mention pipeline-bubbles. Recently proposed Orca [75] also suggests that iteration-level scheduling eliminates bubbles in pipeline scheduling (see Figure 8 in  [75]). However, our experiments show that even with iteration-level scheduling, pipeline bubbles can waste significant GPU cycles with PP (§5.3).
Each micro-batch (or iteration) in LLM inference can require a different amount of compute (and consequently has varying execution time), depending on the composition of prefill and decode tokens in the micro-batch (see Figure 8). We identify three types of bubbles during inference: (1) bubbles like $PB_{1}$ that occur due to the varying number of prefill tokens in two consecutive micro-batches (2) bubbles like $PB_{2}$ that occur due to different compute times of prefill and decode stages when one is followed by the other, and (3) bubbles like $PB_{3}$ that occur due to difference in decode compute times between micro-batches since the attention cost depends on the accumulated context length (size of the KV-cache) and varies across requests. For Falcon-180B, a single prompt of 4k tokens takes $\approx 1150$ ms to execute compared to a decode only iteration with batch size 32 which would take about $\approx 200$ ms to execute. Interleaving of these iteration could result in a bubble of $\approx 950$ ms. These pipeline bubbles are wasted GPU cycles and directly correspond to a loss in serving throughput and increased latency. This problem is aggravated with increase in prompt lengths and batch size, due to longer and more frequent prefill iterations respectively. If we can ensure that each micro-batch performs uniform computation, we can mitigate these pipeline bubbles.
Takeaway-4: There can be a large variance in compute time of LLM iterations depending on composition of prefill- and decode-tokens in the batch. This can lead to significant bubbles when using pipeline-parallelism.
@@ 4 Sarathi-Serve: Design and Implementation
We now discuss the design and implementation of Sarathi-Serve — a system that provides high throughput with predictable tail latency via two key techniques – chunked-prefills and stall-free batching.
@@ 4.1 Chunked-prefills
To tackle this challenge, we present a technique called chunked-prefills which allows computing large prefills in small chunks across several iterations. Chunked-prefills is a prefill splitting mechanism hinged on two key insights. First, as discussed in §3.1, a prefill request with modest sequence length can effectively saturate GPU compute. For example, in Figure 4, prefill throughput starts saturating around sequence length of 512 tokens. Second, in many practical scenarios, input prompts contain several thousand tokens on average (Table 2). This provides an opportunity to break large prefill requests into smaller units of compute which are still large enough to saturate GPU compute. In Sarathi-Serve, we leverage this mechanism to form batches with appropriate number of tokens such that we can utilize the compute potential in decode batches without violating the TBT SLO.
@@
2:Initialize token_budget, $\tau$ $\leftarrow$ compute_token_buget($T_{max}$)
3:Initialize batch_num_tokens, $n_{t}$ $\leftarrow$ 0
@@
7:  if is_prefill_complete($R$) then
@@
10:  if not is_prefill_complete($R$) then
@@ 4.2 Stall-free batching
The Sarathi-Serve scheduler is an iteration-level scheduler that leverages chunked-prefills and coalescing of prefills and decodes to improve throughput while minimizing latency.
[…]
Figure 7 shows the scheduling policy of Sarathi-Serve in action, for the same example used in §3.2. The first iteration is decode-only as there are no prefills to be computed. However, after a new request C enters the system, Sarathi-Serve first splits the prefill of C into two chunks and schedules them in subsequent iterations. At the same time, with stall-free batching, it coalesces the chunked prefills with ongoing decodes of A and B. This way, Sarathi-Serve stalls neither decodes nor prefills unlike existing systems, allowing Sarathi-Serve to be largely free of latency spikes in TBT without compromising throughput. Furthermore, stall-free batching combined with chunked-prefills also ensures uniform compute hybrid batches in most cases, which helps reduce bubbles when using pipeline parallelism, thereby enabling efficient and scalable deployments.
@@ 4.3 Determining Token Budget
During the computation of chunked-prefills, the attention operation for every chunk of a prompt needs to access the KV-cache of all prior chunks of the same prompt. This results in increased memory reads from the GPU HBM even though the computational cost is unchanged. For example, if a prefill sequence is split into $N$ chunks, then the first chunk’s KV-cache is loaded N1N-1 times, the second chunk’s KV-cache is loaded N2N-2 times, and so on. However, we find that even at small chunk sizes attention prefill operation is compute bound operation. In practice, there can be small overhead associated with chunking due to fixed overheads of kernel launch, etc. We present a detailed study of the overheads of chunked-prefills in §5.4.
@@
Another factor that influences the choice of token budget is the tile-quantization effect [13]. GPUs compute matmuls by partitioning the given matrices into tiles and assigning them to different thread blocks for parallel computation. Here, each thread block refers to a group of GPU threads and computes the same number of arithmetic operations. Therefore, matmuls achieve maximum GPU utilization when the matrix dimensions are divisible by the tile size. Otherwise, due to tile-quantization, some thread blocks perform extraneous computation [13]. We observe that tile-quantization can dramatically increase prefill computation time e.g., in some cases, using chunk size of 257 can increase prefill time by 32% compared to that with chunk size 256.
Finally, when using pipeline parallelism the effect of token budget on pipeline bubbles should also be taken into account. Larger chunks lead to higher inter-batch runtime variations that result in pipeline bubbles which results in lower overall system throughput. On the other hand, picking a very small token budget can lead to higher overhead due to lower arithmetic intensity and other fixed overheads.
Therefore, selecting a suitable token budget is a complex decision which depends on the desired TBT SLO, parallelism configuration, and specific hardware properties. We leverage Vidur [28], a LLM inference profiler and simulator to determine the token budget that maximizes system capacity under specific deployment scenario.
4.4 Implementation
We implement Sarathi-Serve on top of the open-source implementation of vLLM [53, 23]. We added support for paged chunk prefill using FlashAttention v2 [38] and FlashInfer [74] kernels. We use FlashAttention backend for all the evaluations in this paper due to its support for wider set of models. We also extend the base vLLM codebase to support various scheduling policies, chunked prefills, pipeline parallelism and an extensive telemetry system. We use NCCL [15] for both pipeline and tensor parallel communication. Source code for the project is available at  https://github.com/microsoft/sarathi-serve.
@@ 5 Evaluation
    What is the maximum load a model replica can serve under specific Service Level Objective (SLO) constraints with different inference serving systems (§5.1) and how does this load vary with varying SLO constraints (§5.2)?
@@
    How does Sarathi-Serve perform under various deployments such as TP and $P$? (§5.3)
@@
    What is the overhead of chunked-prefills? (§5.4.1)
  4. 4.
    What is the effect of each of chunked-prefills and stall-free batching in isolation as opposed to using them in tandem? (§5.4.2)
@@ 5.1 Capacity Evaluation
[…]
@@
We observe that in most scenarios, Orca and vLLM violate the P99 TBT latency SLO before they can reach their maximum serviceable throughput. Thus, we observe relaxing the latency target leads to considerable increase in their model serving capacity. In Sarathi-Serve, one can adjust the chunk size based on the desired SLO. Therefore, we use a strict token budget and split prompts into smaller chunks when operating under strict latency SLO. This reduces system efficiency marginally but allows us to achieve lower tail latency. On the other hand, when the latency constraint is relaxed, we increase the token budget to allow more efficient prefills. We use token budget of 2048 and 512 for all models under the relaxed and strict settings, respectively, except for the LLaMA2-70B relaxed configuration where we use token budget of 1536 to reduce the impact of pipeline bubbles. The system performance can be further enhanced by dynamically varying the token budget based on workload characteristics. We leave this exploration for future work.
@@
We further notice that vLLM significantly outperforms Orca under relaxed setting. The reason for this is two-fold. First, Orca batches prompts for multiple requests together (max sequence length * batch size compared to max sequence length in vLLM), which can lead to even higher tail latency in some cases. Second, vLLM supports a much larger batch size compared to Orca. The lower batch size in Orca is due to the lack of PagedAttention and the large activation memory footprint associated with processing batches with excessively large number of tokens.
@@
Finally, note that the capacity of each system is higher for openchat_sharegpt4 dataset compared to the arxiv_summarization dataset. This is expected because prompts in the arxiv_summarization datasets are much longer - 7059 vs 1730 median tokens as shown in Table 2. The larger prompts makes Orca and vLLM more susceptible to latency violations due to higher processing time of these longer prefills.
5.2 Throughput-Latency Tradeoff
To fully understand the throughput-latency tradeoff in LLM serving systems, we vary the P99 TBT latency SLO and observe the impact on system capacity for vLLM and Sarathi-Serve.  Figure 12 shows the results for Mistral-7B and Yi-34B models with five different SLO values for the openchat_sharegpt4 dataset.
We evaluate vLLM with three different batch sizes in an attempt to navigate the latency-throughput trade-off as prescribed by Yu et al. [75]. The maximum capacity of vLLM gets capped due to generation stalls under stringent TBT SLOs. Notably, the capacity of vLLM remains largely identical for all the three batch size settings. This implies that even though PagedAttention enables large batch sizes with efficient memory management – in practical situations with latency constraints, vLLM cannot leverage the large batch size due to the steep latency-throughput tradeoff made by it’s prefill-prioritizing scheduler.
[…]
5.3 Making Pipeline Parallel Viable
We now show that Sarathi-Serve makes it feasible to efficiently serve LLM inference across commodity networks with efficient pipeline parallelism. For these experiments, we run Falcon-180B over two nodes, each with four A100 GPUs, connected over 100 Gbps Ethernet. We evaluate model capacity under three configurations: vLLM with 8-way TP, vLLM with our pipeline-parallel implementation and Sarathi-Serve with pipeline-parallel. For PP configurations, we do 4-way TP within node and 2-way PP across nodes.
13(a) shows the latency for decode-only batches for Falcon-180B with purely tensor parallel TP-8 deployment compared to a TP-4 PP-2 hybrid parallel configuration. We observe that the median latency for tensor parallelism is $\sim 2$$\times$ higher than pipeline parallelism. This is because TP incurs high communication overhead due to cross-node all-reduces.
[…]
(a) TBT (Falcon-180B).
(b) Capacity (Falcon-180B).
Figure 13: TP scales poorly across nodes. (a) Median TBT for decode-only batches: cross node TP increases median TBT by more than $2\times$ compared to a 4-way TP within node and PP across nodes. (b) Capacity under strict (SLO-S) and relaxed (SLO-R) latency SLOs: Sarathi-Serve increases Falcon-180B’s serving capacity by $4.3\times$ and $3.6\times$ over vLLM’s TP-only and hybrid-parallel configurations under strict SLOs.